OpenStudy (anonymous):

if 3(x+2)-5=15+x what is the value of x

OpenStudy (anonymous):

Expand: 3x+6-5 =15 + x Make x the subject: 3x - x = 15 - 6 + 5 x = ?

OpenStudy (anonymous):

14??

OpenStudy (anonymous):

No, 3x = 14. but x = ? @ericmyer

HanAkoSolo (jamierox4ev3r):

x=8

OpenStudy (anonymous):

thank but i want to learn how to do it not just the answer

OpenStudy (anonymous):

3x - x = 15 - 6 + 5 2x = 14 x = 7

OpenStudy (anonymous):

You would be expanding 3(x+2) first by using the distributive property, then making x the subject of the expression.

HanAkoSolo (jamierox4ev3r):

3x+6=15+x -x -x 2x-1=15 +1 +1 2x=16 /2 /2 x=8

HanAkoSolo (jamierox4ev3r):

this is the process

OpenStudy (anonymous):

how did you get the 1 in 2x-1=15

HanAkoSolo (jamierox4ev3r):

you add one to both sides to get rid of the one on the other side

HanAkoSolo (jamierox4ev3r):

my bad the very fist line is supposed to be 3x+6-5=15+x

HanAkoSolo (jamierox4ev3r):

then, 6-5=1, when you add like terms

OpenStudy (anonymous):

couldn't i say the first line is 6x-5=15+x

HanAkoSolo (jamierox4ev3r):

well i don't know where you get 6x... I will map out the process for you again

HanAkoSolo (jamierox4ev3r):

3(x+2)-5=15+x 3x+6-5=15+x 3x+1=15+x -x -x 2x+1=15 -1 -1 2x=14 /2 /2 x=7 *the answer is different than from before because i caught a small calc. error from before.*

OpenStudy (anonymous):

so were did you get the your 6

HanAkoSolo (jamierox4ev3r):

when you distribute the parentheses, 3*x=3x, and 3*2=6

OpenStudy (anonymous):

awwwwwwww I see

HanAkoSolo (jamierox4ev3r):

okay do you understand?

OpenStudy (anonymous):

yes you are awsome

HanAkoSolo (jamierox4ev3r):

well thats good let me know if you have another problem